Calculate the pH when 11 mL of 440 mM KOH is titrated into 324 mL of 37 mM weak acid with a pKa of 4.98.
Answer #1

millimoles of weak acid = 0.037 x 324 = 11.988

millimoles of KOH added = 0.40 x 11 = 4.4

after KOH added

salt formed = 4.4 millimoles

acid left = 11.988 - 4.4 = 7.588 millimoles

total moles = 11 + 324 = 335 mL

[salt] = 4.4 / 335 = 0.013 M

[acid] = 7.588 / 335 = 0.023 M

pH = pKa + log [salt] / [acid]

pH = 4.98 + log [0.013] / [0.023]

pH = 4.73